How many feet per second are equivalent to 50 miles per hour?
Answer by nabla(475)

5280ft=1 mile
so 50(5280) feet per hour
3600 seconds in 1 hour
50(5280)/3600=73 and 1/3 ft/second

check:
(73.33333)*3600=264000 feet per hour
264000/5280=50 miles per hour.